OpenStudy (anonymous):

please help me with this question 200(2)^t/3=500(1/4)^t/5

OpenStudy (freckles):

\[200 \cdot 2^\frac{t}{3}=500 \cdot (\frac{1}{4})^\frac{t}{5}\]?

OpenStudy (freckles):

\[200 \cdot 2^{\frac{t}{3}}=500(2^{-2})^\frac{t}{5} \\ 200 \cdot 2^\frac{t}{3}=500(2^\frac{-2t}{5}) \\ 200(2^\frac{t}{3})-500(2^{\frac{-2t}{5}})=0 \\ 100 \cdot 2^{\frac{-2t}{5}}(2 ^{\frac{11t}{15}}-5)=0 \] check my math

OpenStudy (freckles):

also this should make it easier for you to solve for t :)

OpenStudy (anonymous):

how about the log on both sides ?

OpenStudy (janu16):

T=30000/7997

OpenStudy (janu16):

@Tin2015

OpenStudy (freckles):

you just need to solve: \[2^{\frac{11t}{15}}=5\] yes do ln( ) on both sides

OpenStudy (anonymous):

Ok thank you freckles :)